# Resolvent estimates for a function of a linear operator

### v1

@@ 1. Introduction
By a Lipschitz curve in $\mathbb{C}$ we mean a curve that can be parametrized as $z=\zeta\cdot(t+if(t))$, $t\in I$, where $I$ is a finite interval, $f\colon I\to\mathbb{R}$ a Lipschitz function and $\zeta$ a complex constant of modulus $1$. Here we will prove the following partial converse of this result.
@@
It is clear that each Lipschitz curve is $1$-admissible, but not vice versa. If the Assouad dimension of a compact set is less than two, then this set is $p$-admissible for some $p<2$. We refer to [3] for details.
Theorem 1.1.
@@
In the Hilbert space case, this is equivalent to $T_{0}$ being similar to a normal operator. An operator $T\in L(X)$ is called spectral if it has a representation $T=T_{0}+N$, where $T_{0}$ is scalar spectral and $N$ is a quasinilpotent operator such that $NE(\delta)=E(\delta)N$ for any Borel set $\delta\subset\mathbb{C}$. We refer to [4] for the basics on spectral operators and to [6] for their simplified definition.
Each scalar spectral operator has first order growth of the resolvent. More generally, if $T$ is scalar and its quasinilpotent part $N$ satisfies $N^{k}=0$ for some $k\in\mathbb{N}$ (i.e., $N$ is nilpotent), then the resolvent of $T$ has $k$th order growth. Therefore, the following result can be seen as a counterpart of the above Theorem 1.1.
@@ 2. Lemmas on analytic functions
Let $N$ be a compact normal operator on a Hilbert space, whose spectrum is contained in ${D(0,1)}^{\text{\,\rm cl}}$ and is so dense that
@@
Let $T=N\oplus J$, where $J$ is the $2\times 2$ Jordan cell corresponding to eigenvalue $3$. Put $f(z)=z^{2}-3z$. Then $f^{\prime}\neq 0$ on $\sigma(T)$ and $T$ fails to have first order resolvent growth near $\sigma(J)=\{3\}$. However, $f(T)$ has first order resolvent growth. Indeed,

### v2

@@ 1. Introduction
By a Lipschitz curve in $\mathbb{C}$ we mean a curve that can be parametrized as $z=\zeta\cdot(t+if(t))$, $t\in I$, where II is a finite interval, $f\colon I\to\mathbb{R}$ a Lipschitz function and $\zeta$ a complex constant of modulus 11. Here we will prove the following partial converse of this result.
@@
It is clear that each Lipschitz curve is 11-admissible, but not vice versa. If the Assouad dimension of a compact set is less than two, then this set is $p$-admissible for some $p<2$. We refer to [3] for details.
Theorem $1.$
@@
In the Hilbert space case, this is equivalent to $T_{0}$ being similar to a normal operator. An operator $T\in L(X)$ is called spectral if it has a representation $T=T_{0}+N$, where $T_{0}$ is scalar spectral and NN is a quasinilpotent operator such that $NE(\delta)=E(\delta)N$ for any Borel set $\delta\subset\mathbb{C}$. We refer to [4] for the basics on spectral operators and to [6] for their simplified definition.
Each scalar spectral operator has first order growth of the resolvent. More generally, if $T$ is scalar and its quasinilpotent part NN satisfies $N^{k}=0$ for some $k\in\mathbb{N}$ (i.e., NN is nilpotent), then the resolvent of $T$ has $k$th order growth. Therefore, the following result can be seen as a counterpart of the above Theorem $1.$
@@ 2. Lemmas on analytic functions
Let NN be a compact normal operator on a Hilbert space, whose spectrum is contained in ${D(0,1)}^{\text{\,\rm cl}}$ and is so dense that
@@
Let $T=N\oplus J$, where $J$ is the $2\times 2$ Jordan cell corresponding to eigenvalue 33. Put $f(z)=z^{2}-3z$. Then $f^{\prime}\neq 0$ on $\sigma(T)$ and $T$ fails to have first order resolvent growth near $\sigma(J)=\{3\}$. However, $f(T)$ has first order resolvent growth. Indeed,